### v1

# Supervised Score-Based Modeling by Gradient Boosting

Abstract
Score-based generative models can effectively learn the distribution of data by estimating the gradient of the distribution. Due to the multi-step denoising characteristic, researchers have recently considered combining score-based generative models with the gradient boosting algorithm, a multi-step supervised learning algorithm, to solve supervised learning tasks. However, existing generative model algorithms are often limited by the stochastic nature of the models and the long inference time, impacting prediction performances. Therefore, we propose a Supervised Score-based Model (SSM), which can be viewed as a gradient boosting algorithm combining score matching. We provide a theoretical analysis of learning and sampling for SSM to balance inference time and prediction accuracy. Via the ablation experiment in selected examples, we demonstrate the outstanding performances of the proposed techniques. Additionally, we compare our model with other probabilistic models, including Natural Gradient Boosting (NGboost), Classification and Regression Diffusion Models (CARD), Diffusion Boosted Trees (DBT), and Bayesian neural network-based models. The experimental results show that our model outperforms existing models in both accuracy and inference time.
Introduction
Generative artificial intelligence (GAI) models aim to create highly realistic simulations or reproductions of data by learning from the characteristics of the original data. Recently, Score-based models (diffusion models) have become the most powerful families of GAI due to their excellent generation performance (Yang et al. 2023; Dhariwal and Nichol 2021). In these models, a series of noises is first added to the original data through a forward diffusion process. The model then learns the added noise and continuously denoises it to generate new data that follows the distribution of the original dataset. At present, many researchers have applied score-based models to generation tasks in many fields, including image generation, audio generation, video generation, etc., all with excellent performance (Rombach et al. 2021; Kong et al. 2020; Ho et al. 2022).
Supervised learning is another important research area in machine learning different from generative learning. In supervised learning, a model is trained on labeled data with the goal of learning a mapping from inputs to outputs, allowing it to predict the output for unseen data accurately (Cunningham, Cord, and Delany 2008). Regression and classification problems are two basic types of supervised learning tasks, which aim to train a neural network for prediction. For regression problems, this prediction result is the target output corresponding to the input data, while for classification problems, it is the accurate label. The learning performance of this data relationship is closely related to the learning ability of the sophisticated structure of neural network models (He et al. 2016). Particularly, the Gradient Boosting Machine (GBM) is a supervised learning model that combines several weak learners into a strong learner (Bentéjac, Csörgő, and Martínez-Muñoz 2021). The sequential weak learners reduce errors from previous iterations to obtain accurate predictions, especially with large and complex datasets. Recently, GBM algorithms such as XGBoost (Chen and Guestrin 2016) and LightGBM (Ke et al. 2017) have demonstrated state-of-the-art performance in tabular competitions on platforms, including Kaggle.
Due to the multi-step denoising characteristic, researchers have recently considered combining score-based generative models with the gradient boosting algorithm to solve supervised learning tasks (Han and Zhou 2024; Beltran-Velez et al. 2024). However, these previous methods are mainly based on diffusion models, whose random denoising process may lead to unstable prediction results and reduce the robustness. Moreover, diffusion models often require longer denoising steps, increasing model inference time. If accelerated algorithms such as DDIM (Song, Meng, and Ermon 2020) are used, the quality of the generated results will usually decrease, resulting in lower prediction accuracy.
Therefore, in this paper, our goal is to propose a score-based gradient boosting model for supervised learning to enable quick and accurate inference. To achieve this goal, we consider the score-based generative models via denoising score matching (Song and Ermon 2019) since its inference process does not have a specific number of denoising steps. Based on this, we propose Supervised Score-based Model (SSM), a score-based model by gradient boosting.
We summarize our main contributions as follows: 1) We utilize noise-free Langevin Dynamics to establish a connection between score-based generative models and GBM algorithms. 2) we present SSM to predict the maximum log-likelihood estimation point, which estimates errors between the target and input data through a score network. 3) For the training process and inference process, we provide a theoretical analysis to balance inference time and prediction accuracy. 4) Experiments on regression and classification tasks show that SSM can achieve better performance than existing methods and significantly shorten the inference time.
Related Work
Score-based generative models have demonstrated powerful learning capabilities for complex data distribution in many fields (Croitoru et al. 2023; Du et al. 2023; Liu et al. 2024). There are generally two categories of such models: diffusion-based and score matching-based models (Yang et al. 2023). Diffusion-based models consist of a forward diffusion process, which adds Gaussian noises into the original data, making it converge to the standard normal distribution, and a denoising process, which removes the added noises from a random variable following normal distribution (Ho, Jain, and Abbeel 2020; Nichol and Dhariwal 2021). The score matching-based model can be regarded as an energy-based model (Song and Kingma 2021), which aims to learn explicit probability distributions of data by a score function. Denoising score matching is a state-of-the-art approach for score estimation, which also estimates the score function from the disturbed dataset and then denoises through Langevin dynamics (Song and Ermon 2019). From a general perspective, These two types of models can also be viewed as the discretization of stochastic differential equations (Song et al. 2020b).
The GBM is a powerful ensemble learning technique that builds models sequentially, with each new model aiming to correct the errors of its predecessors (Friedman 2001). By combining the predictions of multiple weak learners, GBM creates a strong predictive model that can handle complex data patterns and interactions (Natekin and Knoll 2013). XGBoost, LightGBM, and NGBoost are typical decision trees-based gradient boosting techniques that extend the foundational principles of GBM to enhance efficiency and performance. XGBoost and LightGBM focus on computational speed and scalability (Chen and Guestrin 2016; Ke et al. 2017), while NGBoost introduces probabilistic predictions, offering a novel approach to uncertainty estimation in gradient boosting models (Duan et al. 2020).
Recently, via setting input data in supervised learning as conditions, conditional score-based models have been utilized to solve supervised learning problems (Amit et al. 2021; Rahman et al. 2023; Beltran-Velez et al. 2024). (Zimmermann et al. 2021) proposed a score-based generative classifier for classification tasks. This model predicts classification results through maximum likelihood estimation of different label generation distributions and target values instead of direct prediction. In (Han, Zheng, and Zhou 2022), the authors proposed CARD, a score-based model for classification and regression. They emphasized its ability to solve uncertainty in supervised learning and compared its performance with existing Bayesian neural networks (Blundell et al. 2015; Tomczak et al. 2021). The DBT model, a diffusion boosting paradigm, connects the denoising diffusion generative model and GBM via decision trees. Through experiments on real-world regression tasks, this approach demonstrates the potential of integrating GBM with score-based generative models (Han and Zhou 2024).
Background
Gradient Boosting Machine
The GBM is a supervised learning framework that combines several weak learners into strong learners in an iterative way (Friedman 2001). In GBM, a boosted model $F(x)$ is a weighted linear combination of the weak learners, which can be formulated as
$F(x)=F_{0}(x)+\sum_{k=1}^{m}\alpha_{k}h_{k}(x),$ (1)
where $F_{0}(x)$ is the initial prediction, $\alpha_{k}$ is the weight coefficient, and $h_{k}(x)$ represents weak learner. Let $l(y,F(x))$ represent a measure of data error at the observation $(y,x)$ for the differentiable loss function $l$. To train a GBM, the goal is to explore a function $F$ that minimizes the expected loss $L(F(x))=E_{(x,y)}l(y,F(x))$ where the expectation is taken over the input dataset of $(y,x)$. Since the loss function $l$ is differentiable, the optimal solution $F^{*}(x)$ can be represented as
$F^{*}(x)=F_{0}(x)+\sum_{k=1}^{m}\alpha_{k}\cdot(-g_{k}(x)),$ (2)
where $g_{k}(x)=\nabla_{F_{k-1}(x)}L(F_{k-1}(x))$ is the gradient at optimization step $k$. Therefore, weak learners are trained to estimate the negative gradient term to approximate the optimal solution. Given a trained GBM and input, the predicted result can be obtained via the iterative equation Eq. (1).
Langevin Dynamics
Langevin dynamics is a mathematical model that describes the dynamics of molecular systems. For a continuously differentiable probability distribution $p(x)$ where $x\in\mathbb{R}^{d}$, Langevin dynamics uses the score function, i.e., log-likelihood estimate $\nabla_{x}\log p(x)$, iteratively to obtain samples that fit the distribution $p(x)$ (Welling and Teh 2011). Given a certain step size $\epsilon>0$ and any prior distribution $\pi(x)$, the Langevin dynamics can be expressed as
$x_{t}=x_{t-1}+\frac{\epsilon}{2}\nabla_{x}\log p(x_{t-1})+\sqrt{\epsilon}z_{t},$ (3)
where $t\geq 0$, $x_{0}\sim\pi(x)$, and $z_{t}\sim\mathcal{N}(0,1)$. As a Markov chain Monte Carlo (MCMC) technique (Robert, Casella, and Casella 1999), the Langevin equation takes gradient steps based on the score function and also injects Gaussian noise to capture the distribution $p(x)$ instead of the maximum point of the log-likelihood estimate $\log p(x)$. When $\epsilon\rightarrow 0$ and $t\rightarrow\infty$, the distribution of $x_{t}$ will converge to the target distribution $p(x)$ under some regularity conditions (Roberts and Tweedie 1996).
Denoising Score Matching for Score Estimation
Since the Langevin equation in Eq. (3) only relies on the score function $\log p(x)$, estimating the score of the target distribution is a crucial step. Recent studies (Hyvärinen and Dayan 2005; Song et al. 2020a; Vincent 2011) have explored the use of the score network $s_{\theta}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$, a neural network parameterized by $\theta$, to estimate the score function.
Denoising Score Matching
[…]
$J^{\prime}(\theta)=\frac{1}{2}\mathbb{E}_{q(\widetilde{x}|x)p(x)}[\|s_{\theta}(\widetilde{x})-\nabla_{\widetilde{x}}\log q(\widetilde{x}|x)\|_{2}^{2}].$ (4)
When the added noise is small enough, e.g., $q(x)\approx p(x)$, the optimal score network $s_{\theta^{*}}(x)$, where $\theta^{*}=\arg\min_{\theta}J^{\prime}(\theta)$, can represent the score of original data distribution, i.e., $s_{\theta^{*}}(x)=\nabla_{x}\log q(x)\approx\nabla_{x}\log p(x)$ (Vincent 2011). However, denoising score matching may not accurately estimate the score in areas with low data density due to insufficient training data (Song and Ermon 2019). This issue can be solved by adding larger noise, but it will not satisfy the condition that the added noise is sufficiently small, which can lead to another kind of error.
Noise Conditional Score Network
[…]
$l(\theta;\sigma)=\frac{1}{2}\mathbb{E}_{p(x)}\mathbb{E}_{\widetilde{x}\sim\mathcal{N}(x,\sigma^{2}I)}[\|s_{\theta}(\widetilde{x},\sigma)+\frac{\widetilde{x}-x}{\sigma^{2}}\|_{2}^{2}].$ (5)
Then for all noise levels $\{\sigma_{i}\}_{i=1}^{L}$, the loss is,
$\mathcal{L}(\theta;\{\sigma_{i}\}_{i=1}^{L})=\frac{1}{L}\Sigma_{i=1}^{L}\lambda(\sigma_{i})l(\theta;\sigma_{i}),$ (6)
where $\lambda(\sigma_{i})$ is a coefficient depending on $\sigma_{i}$. After being trained, NCSN generates samples using annealed Langevin dynamics (Song and Ermon 2019) where the denoising process occurs level by level according to the noise level.
Supervised Score-based Model via Denoising Score Matching and Gradient Boosting
In supervised learning, the training set $\mathcal{D}$ comprises multiple input-target pairs, i.e., $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, where $x_{i}\in\mathbb{R}^{m}$ and $y_{i}\in\mathbb{R}^{d}$. The model $f:\mathbb{R}^{m}\rightarrow\mathbb{R}^{d}$ trained on this training set aims to predict the response variable $y$ given an input variable $x$, i.e., $f(x)\approx y$. For a regression problem, the response variable $y$ is a continuous variable, whereas it is a categorical variable for classification.
[…]
Accordingly, the objective in our framework is,
$l(\theta;\sigma,x,y)=\frac{1}{2}\mathbb{E}_{p(y)}\mathbb{E}_{\widetilde{y}\sim\mathcal{N}(y,\sigma^{2}I)}[\|s_{\theta}(\widetilde{y},\sigma,x)+\frac{\widetilde{y}-y}{\sigma^{2}}\|_{2}^{2}].$ (7)
Then for all noise levels $\{\sigma_{i}\}_{i=1}^{L}$ and dataset $\mathcal{D}$, the loss is,
$\mathcal{L}(\theta;\{\sigma_{i}\}_{i=1}^{L})=\frac{1}{L}\Sigma_{i=1}^{L}\lambda(\sigma_{i})\mathbb{E}_{(x,y)\in\mathcal{D}}l(\theta;\sigma_{i},x,y).$ (8)
Due to the multi-step denoising characteristic of score-based generative models and GBM models, analyzing the relationship between the two is more helpful in deploying score-based generative models in supervised learning tasks. Additionally, the choice of model parameters greatly affects the performance of the score-based generative model and the GBM model (Bentéjac, Csörgő, and Martínez-Muñoz 2021). As discussed in (Song and Ermon 2020), we need to design many parameters to ensure the effectiveness of training and inference, including (i) the choices of noise scales $\{\sigma_{i}\}_{i=1}^{L}$; (ii) the step size $\epsilon$ in Langevin dynamics; (iii) the inference steps $t$ in Langevin equation. Therefore, in the following, we provide theoretical analysis to ensure the performance of the SSM on regression and classification problems.
Gradient Boosting
Firstly, we consider the connection between SSM and GBM. For a given input, the denoising score match of SSM provides a perturbed distribution in the solution space whose maximum log-likelihood estimation point is the target value. To predict the maximum log-likelihood estimation point, we can set the loss function $l$ in GBM as the negative log-likelihood estimation, i.e., $-log~{}q(F(x))$, where $q(\widetilde{x}|x)$ is the perturbed distribution. Thus, the optimal solution in Eq. (2) will be the maximum log-likelihood estimation point of the perturbed distribution. Particularly, the iterative equation of GBM in Eq. (1) can be converted to a noise-free Langevin equation with the above loss function, i.e.,
$F_{k}(x)=F_{k-1}(x)+\beta_{k}\cdot\nabla log~{}q(F_{k-1}(x)).$ (9)
Moreover, we plan to train one conditional score network with different noise levels as the conditions, similar to the uniform score network in the score-based generative model. The different noise level conditions represent various weak learners in GBM. Specifically, given an input $x_{I}$ and a noise level $\sigma_{j}$, and a trained score network $s_{\theta}(y,\sigma_{j},x_{I})$, the converted noise-free Langevin equation is
$y_{t}=y_{t-1}+\alpha_{j}\cdot s_{\theta}(y_{t-1},\sigma_{j},x_{I}),$ (10)
where $\alpha_{j}=\epsilon\cdot\sigma_{j}^{2}/\sigma_{L}^{2}$ is the step size.
Next, we analyze the convergence results of the noise-free Langevin equation. Assume that we have a well-trained score network that can estimate the gradient accurately. For a prediction pair $(x_{I},y_{I})$, the noise-free Langevin equation is,
$y_{t}=y_{t-1}-r_{L}\cdot(y_{t-1}-y_{I}),$ (11)
where $r_{L}=\epsilon/\sigma_{L}^{2}$ is called the refinement rate. Similar to GBM, it iteratively reduces the difference between $y_{t}$ and the target $y_{I}$, and then lets $y_{t}$ converge to $y_{I}$ as $t\rightarrow\infty$. Simultaneously, the estimation of score network $s_{\theta}(y,\sigma_{j},x_{I})$ is the error estimation weighted by $\frac{1}{\sigma_{i}}$. We call the inference algorithm following Eq. (11) as error refinement.
Technique 1.
(Error Refinement) In the inference stage, the denoising process follows the GBM framework and works as the noise-free Langevin equation (Eq. (10)) without the term in the original Langevin equation (Eq. (3)).
From this perspective, adding a set of noises aims to fully train the score network to obtain more accurate estimates. In (Song and Ermon 2020), the authors proposed a strategy to guide the selection of noise where the score network can be more accurately estimated. This noise level is a geometric progression based on the “three sigma rule of thumb” (Grafarend 2006), which can ensure that the samples from $p_{\sigma_{i}}(x)$ will cover high density regions of $p_{\sigma_{i-1}}(x)$. Besides, for the initial noise scale, the authors suggest choosing $\sigma_{1}$ to be as large as the maximum Euclidean distance between all pairs of training data points, i.e., $\sigma_{1}\geq\max_{y_{i},y_{j}}\|y_{i}-y_{j}\|_{2}$. Score models based on geometric progression have achieved good performance in a variety of tasks (Yang et al. 2023). Therefore, we also adopt a geometric progression noise level and the technique for choosing the initial noise scale.
Refinement steps
It has been observed that a notorious shortcoming of NCSN and other score-based generative models is their relatively long inference time (Cao et al. 2024). In (Song and Ermon 2019), the refinement steps for different levels of noise are usually the same value. However, according to the error refinement equation (Eq. (11)), the final stage is crucial when deploying a hierarchical denoising process. When the last level of noise can be estimated accurately enough, as the denoising step size increases, we can be confident that the predicted value converges around the ground truth. On the contrary, even if the previous levels of noise are estimated accurately, convergence cannot be guaranteed if the last level is not accurate. Hence, the same refinement steps may waste inference time on denoising processes with large levels of noise. Additionally, even though the authors in (Song and Ermon 2020) analyzed how to select a suitable refinement step for different noise levels, it is based on the different inference process from the error refinement following Eq. (10). Therefore, we propose our strategy for step setting.
Give an input-target pair $(x_{I},y_{I})\in\mathcal{D}$, following the design of noise level (Song and Ermon 2020), the samples from $p_{\sigma_{i}}(y_{I})$ will cover high density regions of $p_{\sigma_{i-1}}(y_{I})$. Therefore, in the process of training, there will be more training samples from regions with low noise levels. Additionally, for a perturbed data $\widetilde{y}_{I}$, since the score network $s_{\theta}(\widetilde{y}_{I},\sigma_{i},x_{I})$ estimates $(\widetilde{y}_{I}-y_{I})/\sigma_{i}^{2}$, the estimation based on small noise level is more accurate, when the loss is small enough. Due to the error refinement continuously reducing the error between the current prediction and the ground truth, a simple idea is to select noise levels based on the current error. Based on this, when the error is reduced to high density areas of noise level $\sigma_{i+1}$, we believe the score estimated by the noise level $\sigma_{i+1}$ to be more accurate than $\sigma_{i}$. For the current estimated prediction, we have the following proposition,
Proposition 1.
For an input-target pair $(x_{I},y_{I})\in\mathcal{D}$, let $y_{0}$ denote the initial prediction point. After $t$ steps of denoising, the current estimated prediction $y_{t}$ satisfies,
$y_{t}=(1-r_{L})^{t}\cdot(y_{0}-y_{I})+y_{I},$ (12)
where $r_{L}=\epsilon/\sigma_{L}^{2}$ is the refinement rate.
When $|1-r_{L}|<1$ and $t\rightarrow\infty$, we have $(1-r_{L})^{t}\cdot(y_{0}(\sigma_{L})-y_{I})\rightarrow 0$, i.e., $\lim_{t\rightarrow\infty}y_{t}=y_{I}$. Therefore, to ensure the stability and convergence of the error refinement, we need to demand $|1-r_{L}|<1$, i.e., the refinement rate satisfies $20<r_{L}<$ 2.
Given noise levels $\{\sigma_{i}\}_{i=1}^{L}$ and the initial prediction point $y_{0}(\sigma_{i})$ for each noise level $\sigma_{i}$, let the $t$-th estimated error be $\|y_{t}-y_{I}\|_{\infty}$, i.e., the maximum value of the error in all dimensions. Based on this, we aim to define an end-signal $\beta(\sigma_{i})$. When the error estimated by noise level $\sigma_{i}$ less than this signal $\beta(\sigma_{i})$, we jump into using the next noise level to denoise from the current estimated prediction. We call the jump step as the switch time denoted as $t_{\sigma_{i}}$. Therefore, we have the following,
Proposition 2.
Given a trained score network $s_{\theta}(y_{t},\sigma_{i},x_{I})$, the end-signal $\beta(\sigma_{i})\in\mathbb{R}$ satisfies,
$\begin{split}\sigma_{i}^{2}\cdot\|s_{\theta}(y_{t_{\sigma_{i}}},\sigma_{i},x_{I})\|_{\infty}&<\beta(\sigma_{i})\leq\\ &~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\sigma_{i}^{2}\cdot\|s_{\theta}(y_{t_{\sigma_{i}}-1},\sigma_{i},x_{I})\|_{\infty},\end{split}$ (13)
where $i<L$. Moreover, when $1<i<L$, the switch time $t_{\sigma_{i}}$ satisfies,
$t_{\sigma_{i}}\leq\log_{|1-r_{L}|}(\frac{\beta(\sigma_{i})}{\beta(\sigma_{i-1})}).$ (14)
Specially, when $i=1$, the switch time $t_{\sigma_{i}}$ satisfies,
$t_{\sigma_{1}}\leq\log_{|1-r_{L}|}(\frac{\beta(\sigma_{1})}{\sigma_{1}}).$ (15)
According to Proposition 2, via comparing the value between the current output of the score network $s_{\theta}(y_{t},\sigma_{i},x_{I})$ and end-signal $\beta(\sigma_{i})$, we can then determine when to switch to the next level of noise for denoising. Additionally, Proposition 2 provides an upper bound estimation of switch time $t_{\sigma_{i}}$. This upper bound estimation is relevant to refinement rate $r_{L}$ and end-signal $\beta(\sigma_{i})$. Therefore, via selecting a suitable refinement rate and end-signal, we can effectively limit the number of denoising steps on the specified noise level to shorten the time cost.
For example, the authors chose a geometric sequence with $\sigma_{1}=1$ and $\sigma_{10}=0.01$, denoising step $T=100$, and $\epsilon=2\times 10^{-5}$ for image generation in (Song and Ermon 2019). If we set $\beta(\sigma_{i})=\gamma\cdot\sigma_{i}$, based on Eq. (14), we can approximately compute that $t_{\sigma_{i}}\approx 3\ll 100=T$. Even though the estimation does not consider errors caused by network training, it is still significantly less than the given number of denoising steps which can save a lot of inference time. In summary, we propose that
Technique 2.
(End-signal) By setting an end-signal $\beta(\sigma_{i})$, the denoising process ends based on the estimated error rather than a fixed step size. Additionally, based on the selected refinement rate $r_{L}$ and noise scales $\{\sigma_{i}\}_{i=1}^{L}$, the upper bound of switch time $t_{\sigma_{i}}$ can be estimated when $i<L$.
Refinement rate
As discussed above, the final prediction result depends directly on the last denoising process. It is determined by three parameters: 1) the last noise level $\sigma_{L}$; 2) step size parameter $\epsilon$; 3) the number of the last refinement step. The first two parameters are also the components of the refinement rate, which affects the number of refinement steps of previous noise levels according to Proposition 2. Therefore, we need to specify the relationship between these parameters and the prediction error.
Proposition 3.
Let $e(y_{t},x_{I})$ represent the error between the estimated score and the ground truth score, i.e.,
$e(y_{t},x_{I})=s_{\theta}(y_{t},\sigma_{L},x_{I})+\frac{y_{t}-y_{I}}{\sigma_{L}^{2}}.$
After $t$ times of denoising, the current estimated prediction $y_{t}$ satisfies
$y_{t}=(1-r_{L})^{t}\cdot(y_{0}(\sigma_{L})-y_{I})+y_{I}+\sum_{i=0}^{t-1}(1-r_{L})^{i}\cdot E_{i}(x_{I}),$ (16)
where $E_{i}(x_{I})=\alpha_{L}\cdot e(y_{i},x_{I})$ represents the neural network error.
To analyze the constraints of these parameters, we need the value $y_{t}$ in Eq. (16) to converge as much as possible to ground truth $y_{I}$. However, when $t\rightarrow\infty$, $y_{t}$ cannot converge to $y_{I}$ due to the estimated error $E_{i}(\cdot)$. Therefore, considering the prediction error between current $y_{t}$ and $y_{I}$, we have the following proposition.
Proposition 4.
Assume the last initial denoising point $y_{0}(\sigma_{L})$ follows Technique 2, i.e., $\|y_{0}(\sigma_{L})-y_{I}\|_{\infty}<\beta(\sigma_{L-1})$. Let $E\in\mathbb{R}$ be the upper bound of $E_{i}(x_{I})$ for all $i\geq 0$ and $(x_{I},y_{I})\in\mathcal{D}$, i.e., $\|E_{i}(x_{I})\|_{2}<E$. For an end-signal $\beta(\sigma_{L})$, when
$|1-r_{L}|^{t}\cdot\sqrt{d}\cdot\beta(\sigma_{L-1})+E\cdot\frac{1-|1-r_{L}|^{t}}{r_{L}}<\beta(\sigma_{L}),$ (17)
where $d$ is the number of dimensions of $y_{I}$. we can guarantee that $\|y_{t}-y_{I}\|_{2}<\beta(\sigma_{L})$.
We can use the loss value obtained during training to approximate the error $E$ caused by the network. The end-signal $\beta(\sigma_{L})$ can also be regarded as the error bound we hope to achieve iteratively. The end-signal can be set based on task accuracy requirements. After the end-signal is selected, we can choose all three parameters based on the condition in Eq. (17). Additionally, the first term $|1-r_{L}|^{t}\cdot\sqrt{d}\cdot\beta(\sigma_{L-1})$ decreases as $t$ increases. On the contrary, the second term $E\cdot\frac{1-|1-r_{L}|^{t}}{r_{L}}$ increases as $t$ increases. On the other hand, we can choose a suitable $T$ where the second term has a higher order than the first term, i.e., $(|1-r_{L}|^{T}\cdot\sqrt{d}\cdot\beta(\sigma_{L-1}))/(E\cdot\frac{1-|1-r_{L}|^{T}}{r_{L}})\ll 1$. When inference time $t\geq T$, we can ensure that the error of the current prediction result is dominated by the network training error. This approach can balance the error caused by the neural network and the error caused by iteration as much as possible. In this case, $t\geq\log_{|1-r_{L}|}\frac{E}{\sqrt{d}\cdot\beta(\sigma_{L-1})\cdot r_{L}+E}$. In summary,
Technique 3.
(Refinement rate) By setting an end-signal $\beta(\sigma_{L})$, choose suitable $r_{L}$ and $T$ in the sampling process following Eq. (17). Additionally, these parameters can make the second term have a higher order of magnitude than the first term in Eq. (17).
Weigh of Loss
The choice of coefficient $\lambda(\cdot)$ in Eq (8) can affect how well the network is trained for different noise levels. In (Song and Ermon 2019), the authors aimed to let the value of $\lambda(\sigma_{i})l(\theta;\sigma_{i})$ in Eq (6) at the same order of magnitude for all $\{\sigma_{i}\}_{i=1}^{L}$. To achieve this, they chose $\lambda(\sigma_{i})$ as $\sigma_{i}^{2}$. Setting the same magnitude, the loss of the score network learns to a similar degree for each level of noise in each back-propagation and gradient descent.
When applying Techniques 2 and 3, the number of denoising steps for different noises is different, and the last step of denoising is the most important. Based on this, to obtain a better estimation of the score function for the last noise level, an intuitive idea is to enhance the training weight in the loss function (Caruana 1997). Therefore, we choose a different coefficient $\lambda(\sigma_{i})$ in Eq (8), specifically taken as $\sigma_{i}^{k}$ and $k<2$. Since $\{\sigma_{i}\}_{i=1}^{L}$ is a geometric progression and $\sigma_{i}^{2}l(\theta;\sigma_{i})\propto 1$, $\sigma_{i}^{k}l(\theta;\sigma_{i})\propto\sigma_{i}^{k-2}$ is also a geometric progression getting the maximum value when $i=L$, i.e., the last noise level. Based on this, we propose a coefficient selection strategy.
Technique 4.
(Weigh of Loss) Choosing suitable coefficients $\lambda(\sigma_{i})$ in the loss (Eq. (8)) increases the magnitude of the value $\lambda(\sigma_{i})l(\theta;\sigma_{i},x,y)$ as $i$ increases, such as $\lambda(\sigma_{i})=\sigma_{i}$.
By employing Techniques 14 together, the complete training and inference procedure are shown as Alg. 1 and Alg. 2.
Algorithm 1 Training
0:  $\{\sigma_{i}\}_{i=1}^{L}$, training set $\mathcal{D}$, loss coefficient $\lambda(\sigma_{i})$
1:  repeat
2:     choose $(x,y)\in\mathcal{D}$, $\sigma\in\{\sigma_{i}\}_{i=1}^{L}$, and $\widetilde{y}\sim\mathcal{N}(y,\sigma)$
3:     Take gradient descent step on
4:     $\nabla_{\theta}\frac{1}{2}\lambda(\sigma_{i})\|s_{\theta}(\widetilde{y},\sigma,x)+\frac{\widetilde{y}-y}{\sigma^{2}}\|_{2}^{2}$
5:  until
6:  converged
Algorithm 2 Inference
0:  $\{\sigma_{i}\}_{i=1}^{L},\varepsilon,T,\{\beta(\sigma_{i})\}_{i=1}^{L-1},x_{I}$
1:  Initialize $y_{0}$
2:  for $i\leftarrow 1$ to $L-1$ do
3:     $\alpha_{i}\leftarrow\epsilon\cdot\frac{\sigma_{i}^{2}}{\sigma_{L}^{2}}$
4:     repeat
5:        $y_{t}\leftarrow y_{t-1}+\alpha_{i}s_{\theta}(y_{t-1},\sigma_{i},x_{I})$
6:     until $\sigma_{i}^{2}\cdot\|s_{\theta}(y_{t},\sigma_{i},x_{I})\|_{\infty}<\beta(\sigma_{i})$
7:     $y_{0}\leftarrow y_{T}$
8:  end for
9:  for $t\leftarrow 1$ to $T$ do
10:     $y_{t}\leftarrow y_{t-1}+\epsilon s_{\theta}(y_{t-1},\sigma_{L},x_{I})$
11:  end for
12:  return $y_{T}$
Figure 1: The scatter plots for toy examples. From left to right: linear regression, quadratic regression, log-log linear regression, log-log cubic regression, and sinusoidal regression. The green and blue points represent the true values and the prediction results generated by 1000 samples, respectively.
Task Technique RMSE $\downarrow$
$\lambda(\sigma_{i})=\sigma_{i}$ (L1) $\lambda(\sigma_{i})=\sigma_{i}^{2}$ (L2)
w/o noise, fast w/ noise, fast w/ noise w/o noise w/ noise, fast w/o noise w/o noise, fast w/ noise
Linear 0.0667 $0.0668$ $0.1767$ $0.0773$ $0.0685$ $0.0822$ $0.0684$ $0.1764$
Quadratic $0.1197$ $0.1198$ $0.1103$ $0.1296$ $0.0951$ $0.1026$ 0.0950 $0.1139$
Log-Log Linear $0.0783$ $0.0781$ $0.0857$ 0.0654 $0.0793$ $0.0885$ $0.0794$ $0.0699$
Log-Log Cubic $3.2704$ $3.2213$ $3.3903$ $3.9410$ 1.9919 $2.9880$ $2.0050$ $3.2718$
Sinusoidal $0.0096$ $0.0097$ $0.0098$ $0.0094$ $0.0035$ $0.0033$ 0.0033 $0.0034$
Table 1: RMSE on different ablation settings on toy examples. W/ and w/o noise represent whether the noise term is included in Langvin dynamics; fast denotes the use of Technique 2; L1 and L2 represent the degree of noise levels of 1 and 2, respectively.
Task Technique Inference Time (s) $\downarrow$
$\lambda(\sigma_{i})=\sigma_{i}$ (L1) $\lambda(\sigma_{i})=\sigma_{i}^{2}$ (L2)
w/o noise, fast w/ noise, fast w/ noise w/o noise w/ noise, fast w/o noise w/o noise, fast w/ noise
Linear $0.0385$ 0.0354 $0.1404$ $0.1114$ $0.0376$ $0.1106$ $0.0366$ $0.1404$
Quadratic 0.2743 $0.2774$ $1.0957$ $0.9622$ $0.3262$ $0.9631$ $0.3252$ $1.0909$
Log-Log Linear $0.4151$ $0.4176$ $1.1151$ $0.9771$ $0.3425$ $0.9647$ 0.3415 $1.0930$
Log-Log Cubic $0.1311$ $0.1319$ $0.5741$ $0.5015$ 0.0987 $0.4996$ $0.0998$ $0.5843$
Sinusoidal $0.1774$ 0.1766 $1.0863$ $0.9545$ $0.2103$ $0.9590$ $0.2066$ $1.1023$
Table 2: Inference time on different ablation settings on toy examples. W/ and w/o noise represent whether the noise term is included in Langvin dynamics; fast denotes the use of Technique 2; L1 and L2 represent the degree of noise levels of 1 and 2, respectively.
Dataset RMSE $\downarrow$
NGboost MC Dropout Deep Ensembles GCDS CARD DBT SSM (L1) SSM (L2)
Boston 2.94 $\pm$ 0.53 3.06 $\pm$ 0.96 3.17 $\pm$ 1.05 2.75 $\pm$ 0.58 2.61 $\pm$ 0.63 2.73 $\pm$ 0.62 2.06 $\pm$ 0.19 2.04 $\pm$ 0.14
Concrete 5.06 $\pm$ 0.61 5.09 $\pm$ 0.60 4.91 $\pm$ 0.47 5.39 $\pm$ 0.55 4.77 $\pm$ 0.46 4.56 $\pm$ 0.50 5.88 $\pm$ 1.00 3.20 $\pm$ 1.00
Energy 0.46 $\pm$ 0.06 1.70 $\pm$ 0.22 2.02 $\pm$ 0.32 0.64 $\pm$ 0.09 0.52 $\pm$ 0.07 0.52 $\pm$ 0.07 0.36 $\pm$ 0.09 0.31 $\pm$ 0.02
Kin8nm 16.03 $\pm$ 0.04 7.10 $\pm$ 0.26 8.65 $\pm$ 0.47 8.88 $\pm$ 0.42 6.32 $\pm$ 0.18 7.04 $\pm$ 0.23 8.30 $\pm$ 0.10 4.66 $\pm$ 0.30
Naval 0.13 $\pm$ 0.00 0.08 $\pm$ 0.03 0.09 $\pm$ 0.01 0.14 $\pm$ 0.05 0.02 $\pm$ 0.00 0.07 $\pm$ 0.01 0.10 $\pm$ 0.00 0.01 $\pm$ 0.00
Power 3.79 $\pm$ 0.18 4.04 $\pm$ 0.14 4.02 $\pm$ 0.15 4.11 $\pm$ 0.16 3.93 $\pm$ 0.17 3.95 $\pm$ 0.16 4.25 $\pm$ 0.01 3.67 $\pm$ 0.33
Protein 4.33 $\pm$ 0.03 4.16 $\pm$ 0.12 4.45 $\pm$ 0.02 4.50 $\pm$ 0.02 3.73 $\pm$ 0.01 3.81 $\pm$ 0.04 5.20 $\pm$ 0.01 3.95 $\pm$ 0.43
Wine 0.63 $\pm$ 0.04 0.62 $\pm$ 0.04 0.63 $\pm$ 0.04 0.66 $\pm$ 0.04 0.63 $\pm$ 0.04 0.61$\pm$ 0.04 0.68 $\pm$ 0.01 0.46 $\pm$ 0.21
Yacht 0.50 $\pm$ 0.20 0.84 $\pm$ 0.27 1.19 $\pm$ 0.49 0.79 $\pm$ 0.26 0.65 $\pm$ 0.25 1.08 $\pm$ 0.39 0.56 $\pm$ 0.16 0.37 $\pm$ 0.21
Year 8.94 $\pm$ NA 8.77 $\pm$ NA 8.79 $\pm$ NA 9.20 $\pm$ NA 8.70 $\pm$ NA 8.81 $\pm$ NA 10.14 $\pm$ NA 10.04 $\pm$ NA
# best 0 0 0 0 2 0 0 8
Table 3: RMSE results for UCI regression tasks. For Kin8nm and Naval dataset, the results are multiplied by 100 to match the scale of others.
Model CARD SSM (L1) SSM (L2)
CIFAR Accuracy 90.93 $\pm$ 0.02 90.91 $\pm$ 0.00 90.99 $\pm$ 0.00
-10 Time 50.30 $\pm$ 0.30 19.63 $\pm$ 1.21 15.43 $\pm$ 1.12
CIFAR Accuracy 71.42 $\pm$ 0.01 71.51 $\pm$ 0.00 71.38 $\pm$ 0.00
-100 Time 90.58 $\pm$ 0.12 32.30 $\pm$ 0.60 27.16 $\pm$ 0.19
Table 4: Accuracy (in %) and inference time for one batch (in second (s)) by CARD and SSM on CIFAR-10 and CIFAR-100.
Experiment
Toy examples
To demonstrate the effectiveness of SSM and all improved Techniques, we first perform experiments on 5 selected toy examples (linear regression, quadratic regression, log-log linear regression, log-log cubic regression, and sinusoidal regression) proposed in (Han, Zheng, and Zhou 2022). We add unbiased normal distribution noise to the original models to obtain noisy data for training, increasing model complexity. We aim to predict accurate regression results for pure input data in the inference process. To evaluate the impact of various technologies, we have designed an ablation experiment that considers three variables: the use of original Langevin dynamics, the implementation of the fast sampling technique as described in Technique 2, and the application of different coefficients $\lambda(\sigma_{i})$ in the loss function, where L1 and L2 represent the degree of noise levels of 1 and 2, respectively. The experiment results with the Root Mean Squared Error (RMSE) and inference time about different model settings are shown in Table 1 and 2, respectively. The scatter plots for 5 examples of both true and generated data following the first setting, i.e., L1, w/o noise fast, are shown in Figure 1. We observe that while perturbed data is used, SSM can still fit the regression equation of the original data. The proposed fast sampling Technique 2 significantly reduces the inference time for all tasks. Additionally, since the fast sampling technique ensures the error bound of the current estimated state, it can even achieve the best performance on tasks except for log-log linear regression. Using the noise-free Langevin equation, i.e., w/o noise, performs better on most tasks, indicating that the selection of different coefficients also affects the performance of the model.
Regression
We further evaluate our model on 10 UCI regression tasks (Dua and Graff 2017). We employ the same experimental settings as those used in the CARD model (Han, Zheng, and Zhou 2022). We compare our model with NGboost, CARD, DBT, and several Bayesian neural network frameworks considered in (Han, Zheng, and Zhou 2022), including MC Dropout (Gal and Ghahramani 2016), Deep Ensembles (Lakshminarayanan, Pritzel, and Blundell 2017), and GCDS (Zhou et al. 2023). The experiment results with RMSE are shown in Table 3. It is noticed that SSM outperforms existing approaches in RMSE. SSM (L2) obtains the best results in 8 out of 10 datasets. Although the RMSE of SSM (L1) is larger, it has a smaller standard deviation, which means the network is more stable. In addition, probabilistic models such as CARD predict the probability of distribution, while SSM predicts the points directly. This also leads to a lower standard deviation of SSM on most tasks.
Classification
For classification tasks, we compare our model with CARD on CIFAR-10 and CIFAR-100, focusing on both accuracy and the inference time (Krizhevsky 2009). Table 4 shows the comparison results of accuracy for CIFAR-10 and CIFAR-100 classification. We observe that SSM can achieve excellent performance for prediction accuracy with lower variance. This shows that our method has more stable prediction results when selecting different random number seeds. Additionally, SSM can significantly shorten the inference time while ensuring performance compared with CARD, which has the same network structure.
Conclusion
In this paper, we proposed Supervised Score-based Model (SSM), score-based gradient boosting model using denoising score matching for supervised learning. First, we analyzed the connection between SSM and GBM and showed that the iterative equation of GBM can be converted to a noise-free Langevin equation. Then, we provided a theoretical analysis on learning and sampling for SSM to balance inference time and prediction accuracy with noise-free Langevin sampling. Furthermore, we showed the effectiveness of the proposed techniques on selected toy examples. Lastly, we compared SSM with existing models, including NGboost, CARD, DBT, and Bayesian neural network-based models, for several regression and classification tasks. The experimental results show that our model outperforms in both accuracy and inference time.